# Vector valued estimates for matrix weighted maximal operators and product $\mathrm{BMO}$

### v1

@@ Abstract
We consider maximal operators acting on vector valued functions, that is functions taking values on $\mathbb{C}^{d},$ that incorporate matrix weights in their definitions. We show vector valued estimates, in the sense of Fefferman–Stein inequalities, for such operators. These are proven using an extrapolation result for convex body valued functions due to Bownik and Cruz-Uribe. Finally, we show an $\mathrm{H}^{1}$-$\mathrm{BMO}$ duality for matrix valued functions and we apply the previous vector valued estimates to show upper bounds for biparameter paraproducts.
@@ 1.1 Main results
We deduce Theorem 1 from a general principle for establishing matrix weighted bounds for vector valued extensions of operators acting on convex body valued functions, see Theorem 5 below. Our method for deducing such bounds is inspired from [Book_Extrapolation]: we use an analog of the Rubio de Francia extrapolation theorem for matrix weights proved in [Cruz_Uribe_Bownik_Extrapolation], coupled with a trick of interpreting vector valued extensions of operators as operators whose values are convex body valued functions. It is worth noting that using the exact same method, the very recent limited range extrapolation theorem for matrix weights proved in [Kakaroumpas_Nguyen_Vardakis_2024] yields similar bounds as in Theorem 5 that are valid only for a limited range of exponents.
@@ Acknowledgements
The authors are grateful to S. Treil for suggesting the interpretation of the $q$-function as an operator whose values are convex body valued functions for any $1<q<\infty,$ inspiring the interpretation of vector valued extensions of operators as operators whose values are convex body valued functions used in this paper.
@@ 2.7 Convex body valued functions
Here we consider functions taking values in the collection of closed bounded symmetric convex sets of $\mathbb{C}^{d}.$ Take into account that by symmetric here we mean complex symmetric. That is, a set $A\subseteq\mathbb{C}^{d}$ is complex symmetric (or just symmetric in this article) if for every $u\in A$ and every $\lambda\in\mathbb{C}$ with $|\lambda|=1$ it is also the case that $\lambda u\in A.$ Of course, if the set $A$ is convex in addition to symmetric, it will also be the case that, for every $u\in A$ and every $\lambda\in\mathbb{C}$ with $|\lambda|\leq 1,$ also $\lambda u\in A.$ We will denote the set of closed subsets of $\mathbb{C}^{d}$ by $\mathcal{K}(\mathbb{C}^{d}),$ or just $\mathcal{K}$ when the dimension of the ambient space is clear by the context. In addition, we define a convex body to be a closed bounded convex and symmetric subset of $\mathbb{C}^{d}.$ We will use the symbol $\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{d})$ to denote the set of convex bodies on $\mathbb{C}^{d}$ and, whenever the dimension of the ambient space is unambiguous, simply by $\mathcal{K}_{\mathrm{bcs}}.$ We focus now on functions $F\colon\mathbb{R}^{n}\rightarrow\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{d})$ and we gather the definitions and basic properties that we will need for our results. Some of these definitions can be found in more general forms in the texts that we cite, we will restrict though to the cases that are of interest to us to avoid an excess of concepts. For an introduction to such functions, their properties and how to define their integrals see [Cruz_Uribe_Bownik_Extrapolation, Cruz_Uribe_Extrapolation] and for a detailed exposition see [Aubin_2009].
@@
$F(x)=\overline{\{f_{k}(x)\colon k\geq 1\}}$
for every $x\in\mathbb{R}^{n}$ (see [Aubin_2009, Theorem 8.1.4]). The functions that form such sequences will be called selection functions for $F.$ In addition, the set of all selection functions for $F$ is denoted by $S^{0}(F).$ Observe that this is the same as the set of all measurable functions $f$ such that $f(x)\in F(x)$ for every $x\in\mathbb{R}^{n}.$ When one is interested in the norm of $F,$ it is possible to restrict to selection functions, as the following lemma shows (see [Cruz_Uribe_Bownik_Extrapolation, Lemma 3.9]).
@@
$A_{Q}F(x)\coloneq\frac{\chi_{Q}(x)}{|Q|}\int_{\mathbb{R}^{n}}F(y)\,\,\mathrm{dm}(y),$
@@ 4.1 $\mathrm{H}^{1}$-BMO duality
$\|B\|_{\emph{BMO}_{\emph{prod},\bm{\mathcal{D}}}(U,V,p)}\lesssim_{n,m,p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{2}\|\ell\|.$
@@
By testing the strong dyadic Christ–Goldberg maximal operator $M_{\bm{\mathcal{D}},V^{\prime}}$ with respect to the $(d\times d)$ matrix $\bm{\mathcal{D}}$-dyadic biparameter matrix $A_{p^{\prime}}$ weight $V^{\prime}$ on the function $V^{1/p}\mathbf{1}_{\Omega}e_{k}$ for each $k=1,\ldots,d$ (where we recall that $\{e_{1},\ldots,e_{d}\}$ is the standard basis of $\mathbb{C}^{d}$), and because this operator is bounded (4.2), we obtain
$\|N_{\Omega}\|_{L^{p^{\prime}}(\mathbb{R}^{2})}\lesssim_{n,m,p,d}[V^{\prime}]_{A_{p^{\prime}},\bm{\mathcal{D}}}^{2/(p^{\prime}-1)}|\Omega|^{1/p^{\prime}}\sim_{p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{2}|\Omega|^{1/p^{\prime}}.$
@@
$\lesssim_{n,m,p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{2}|\Omega|^{1/p^{\prime}}\|A\|_{\mathrm{L}^{p}(\mathbb{R}^{n+m};\mathrm{M}_{d}(\mathbb{C}))},$
@@
It is obvious that the proofs of Propositions 8, 9 work also in the one-parameter setting.

### v2

@@ Abstract
We consider maximal operators acting on vector valued functions, that is functions taking values on $\mathbb{C}^{d},$ that incorporate matrix weights in their definitions. We show vector valued estimates, in the sense of Fefferman–Stein inequalities, for such operators. These are proven using an extrapolation result for convex body valued functions due to Bownik and Cruz-Uribe. Finally, we show an $\mathrm{H}^{1}$-$\mathrm{BMO}$ duality for matrix valued functions and we apply the previous vector valued estimates to show upper bounds for biparameter paraproducts. For the reader’s convenience, we include an appendix explaining how to adapt the extrapolation for real convex body valued functions of Bownik and Cruz-Uribe to the setting of complex convex body valued functions that we treat.
@@ 1.1 Main results
We deduce Theorem 1 from a general principle for establishing matrix weighted bounds for vector valued extensions of operators acting on convex body valued functions, see Theorem 5 below. Our method for deducing such bounds is inspired from [Book_Extrapolation]: we use an analog of the Rubio de Francia extrapolation theorem for matrix weights proved in [Cruz_Uribe_Bownik_Extrapolation], coupled with a trick of interpreting vector valued extensions of operators as operators whose values are convex body valued functions. It is worth noting that using the exact same method, the recent limited range extrapolation theorem for matrix weights proved in [Kakaroumpas_Nguyen_Vardakis_2024] yields similar bounds as in Theorem 5 that are valid only for a limited range of exponents. Note that the extrapolation for matrix weights and the other methods and techniques from [Cruz_Uribe_Bownik_Extrapolation] (which also belong to the foundations of the work in [Kakaroumpas_Nguyen_Vardakis_2024]) concern the setting of real convex body valued functions. An extension to complex convex body valued functions presents nevertheless no difficulties thanks to previous work in [DKPS2024]. For the reader’s convenience we supply details in the appendix.
@@ Acknowledgements
The authors are grateful to S. Treil for suggesting the interpretation of the $q$-function as an operator whose values are convex body valued functions for any $1<q<\infty,$ inspiring the interpretation of vector valued extensions of operators as operators whose values are convex body valued functions used in this paper. In addition, the authors are indebted to both Marcin Bownik and David Cruz-Uribe for carefully reading this article and for providing valuable feedback that improved the clarity and readability of this work.
@@ 2.7 Convex body valued functions
Here we consider functions taking values in the collection of closed bounded symmetric convex sets of $\mathbb{C}^{d}.$ Take into account that by symmetric here we mean complex symmetric. That is, a set $A\subseteq\mathbb{C}^{d}$ is complex symmetric (or just symmetric in this article) if for every $u\in A$ and every $\lambda\in\mathbb{C}$ with $|\lambda|=1$ it is also the case that $\lambda u\in A.$ Of course, if the set $A$ is convex in addition to symmetric, it will also be the case that, for every $u\in A$ and every $\lambda\in\mathbb{C}$ with $|\lambda|\leq 1,$ also $\lambda u\in A.$ In other words, the “symmetric convex sets” as we have defined them are precisely the balanced convex sets. We will denote the set of closed subsets of $\mathbb{C}^{d}$ by $\mathcal{K}(\mathbb{C}^{d}),$ or just $\mathcal{K}$ when the dimension of the ambient space is clear by the context. In addition, we define a convex body to be a closed bounded convex and symmetric subset of $\mathbb{C}^{d}.$ We will use the symbol $\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{d})$ to denote the set of convex bodies on $\mathbb{C}^{d}$ and, whenever the dimension of the ambient space is unambiguous, simply by $\mathcal{K}_{\mathrm{bcs}}.$ We focus now on functions $F\colon\mathbb{R}^{n}\rightarrow\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{d})$ and we gather the definitions and basic properties that we will need for our results. Some of these definitions can be found in more general forms in the texts that we cite, we will restrict though to the cases that are of interest to us to avoid an excess of concepts. For an introduction to such functions, their properties and how to define their integrals see [Cruz_Uribe_Bownik_Extrapolation, Cruz_Uribe_Extrapolation] and for a detailed exposition see [Aubin_2009].
@@
$F(x)=\overline{\{f_{k}(x)\colon k\geq 1\}}$ (2.5)
for every $x\in\mathbb{R}^{n}$ (see [Aubin_2009, Theorem 8.1.4]). The functions that form such sequences will be called selection functions for $F.$ In addition, the set of all selection functions for $F$ is denoted by $S^{0}(F).$ Observe that this is the same as the set of all measurable functions $f$ such that $f(x)\in F(x)$ for every $x\in\mathbb{R}^{n}.$ Here we restrict to functions taking values on $\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{d})$ due to the applications that we consider later. Nonetheless, the previous definitions and concepts apply verbatim to functions taking values on $\mathcal{K}(\mathbb{C}^{d}).$ When one is interested in the norm of $F,$ it is possible to restrict to selection functions, as the following lemma shows (see [Cruz_Uribe_Bownik_Extrapolation, Lemma 3.9]). We include its proof for completeness, although the arguments of Bownik and Cruz-Uribe are equally valid for complex convex bodies in this case.
@@
Proof.
Through this proof, we identify $\mathbb{C}^{d}$ with $\mathbb{R}^{2d}$ in the usual way. Take the sequence $\{f_{k}\}$ of selection functions satisfying (2.5). Then, the function given by
$g_{0}(x)=\sup\{|v|\colon v\in F(x)\}=\sup\{|f_{k}(x)|\colon k\geq 1\}$
is also measurable. This allows us to define the function
$F_{0}(x)=\{v\in F(x)\colon|v|=g_{0}(x)\}=F(x)\cap\mathbf{{S}}_{\mathbb{C}^{{d}}}$
taking values on $\mathcal{K}(\mathbb{C}^{d}),$ where $\mathbf{{S}}_{\mathbb{C}^{{d}}}=\{v\in\mathbb{C}^{d}\colon|v|=1\}$ denotes the complex $(d-1)$-dimensional sphere. Since both $F$ and $g_{0}\mathbf{{S}}_{\mathbb{C}^{{d}}}$ are measurable, so is $F_{0}$ (see [Aubin_2009, Theorems 8.2.2 and 8.2.4]).
It is only left to choose $v(x)\in F_{0}(x)$ in a measurable way. This is done by choosing maximal vectors $v$ in every other real coordinate iteratively (in the $\mathbb{R}^{2d}$ sense). Let $P_{1}$ be the (continuous) projection onto the first complex coordinate and define
$g_{1}(x)=\sup_{k}|P_{1}(f_{k}(x))|.$
The function $g_{1}\colon\Omega\rightarrow[0,\infty)$ is then measurable. Next, use this function and the symmetry of $F(x)$ to define $F_{1}\colon\Omega\rightarrow\mathcal{K}(\mathbb{C}^{d})$ as
$F_{1}(x)=\{v\in F_{0}(x)\colon P_{1}(v)=g_{1}(x)\}=F_{0}(x)\cap(\{g_{1}(x)\}\times\mathbb{C}^{d-1}).$
As before, $F_{1}$ is also a measurable function. Assume that we have defined $F_{j}$ measurable and taking values on $\mathcal{K}(\mathbb{C}^{d}).$ Define $F_{j+1}\colon\Omega\rightarrow\mathcal{K}(\mathbb{C}^{d})$ by taking the set of points of $F_{j}(x)$ with maximal modulus of the $j+1$ complex coordinate, positive real part and zero imaginary part for the same coordinate. Also $F_{j+1}$ will be measurable because of the same reasons as before. Eventually, we get the measurable function $F_{d}\colon\Omega\rightarrow\mathcal{K}(\mathbb{C}^{d})$ and $F_{d}(x)$ must be a singleton for every $x\in\Omega,$ this is $F_{d}(x)=\{v(x)\}$ for some $v(x)\in\mathbb{C}^{d},$ because of the maximality of the modulus of each complex coordinate and the restriction of having each complex coordinate lying on the positive real axis. ∎
@@
$A_{Q}F(x)\coloneq\frac{\mathbf{1}_{Q}(x)}{|Q|}\int_{Q}F(y)\,\,\mathrm{dm}(y),$
@@ 4.1 $\mathrm{H}^{1}$-BMO duality
$\|B\|_{\emph{BMO}_{\emph{prod},\bm{\mathcal{D}}}(U,V,p)}\lesssim_{n,m,p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{2+1/p}\|\ell\|.$
@@
We define now functions $\widetilde{N}_{\Omega}^{(k)}$ by applying the strong dyadic Christ–Goldberg maximal operator $M_{\bm{\mathcal{D}},V^{\prime}}$ with respect to the biparameter $(d\times d)$ matrix valued $\bm{\mathcal{D}}$-dyadic $A_{p^{\prime}}$ weight $V^{\prime}$ on the function $V^{1/p}\mathbf{1}_{\Omega}e_{k}$ for each $k=1,\ldots,d$ (where we recall that $\{e_{1},\ldots,e_{d}\}$ is the standard basis of $\mathbb{C}^{d}$). That is, we define
$\widetilde{N}_{\Omega}^{(k)}(x)\coloneq\sup_{R\in\bm{\mathcal{D}}}\langle|V(x)^{-1/p}V^{1/p}\mathbf{1}_{\Omega}e_{k}|\rangle_{R}\mathbf{1}_{R}(x).$
If we consider $\widetilde{N}_{\Omega}\coloneq\sup_{k=1,\ldots,d}\widetilde{N}_{\Omega}^{(k)},$ this satisfies
$\widetilde{N}_{\Omega}(x)\gtrsim_{n,m,d,p}[V]_{A_{p},\bm{\mathcal{D}}}^{-1/p}N_{\Omega}(x).$
Indeed, just observe that
$N_{\Omega}(x)$ $=\sup_{R\in\bm{\mathcal{D}}(\Omega)}|V(x)^{-1/p}\mathcal{V}_{R}|\mathbf{1}_{R}(x)=\sup_{R}|\mathcal{V}_{R}V(x)^{-1/p}|\mathbf{1}_{R}(x)$
$\sim_{d}\sup_{R\in\bm{\mathcal{D}}(\Omega)}\langle|V^{1/p}V(x)^{-1/p}|^{p}\rangle^{1/p}\mathbf{1}_{R}(x)$
$\lesssim_{n,m,d,p}[V]_{A_{p},\bm{\mathcal{D}}}^{1/p}\sup_{R\in\bm{\mathcal{D}}(\Omega)}\langle|V^{1/p}V(x)^{-1/p}|\rangle_{R}\mathbf{1}_{R}(x)$
$=[V]_{A_{p},\bm{\mathcal{D}}}^{1/p}\sup_{R\in\bm{\mathcal{D}}(\Omega)}\langle|V(x)^{-1/p}V^{1/p}|\rangle_{R}\mathbf{1}_{R}(x),$
because the scalar weight $|V^{1/p}M|^{p}$ is uniformly in the biparameter $\bm{\mathcal{D}}$-dyadic Muckenhoupt $A_{p}$ class for every positive definite matrix $M$ (see for example [DKPS2024, Lemma 3.4]). Next, using the comparability between the matrix norm and the supremum of norms of matrix columns, we get
$N_{\Omega}(x)$ $\lesssim_{n,m,d,p}[V]_{A_{p},\bm{\mathcal{D}}}^{1/p}\sup_{R\in\bm{\mathcal{D}}(\Omega)}\sup_{k=1,\ldots,d}\langle|V(x)^{-1/p}V^{1/p}e_{k}|\rangle_{R}\mathbf{1}_{R}(x)$
$\leq[V]_{A_{p},\bm{\mathcal{D}}}^{1/p}\widetilde{N}_{\Omega}(x).$
Finally, using the boundedness (4.2) of the strong dyadic Christ–Goldberg maximal operator, we obtain
$\|N_{\Omega}\|_{L^{p^{\prime}}(\mathbb{R}^{2})}\lesssim_{n,m,p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{1/p}[V^{\prime}]_{A_{p^{\prime}},\bm{\mathcal{D}}}^{2/(p^{\prime}-1)}|\Omega|^{1/p^{\prime}}\sim_{p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{2+1/p}|\Omega|^{1/p^{\prime}}.$
@@
$\lesssim_{n,m,p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{2+1/p}|\Omega|^{1/p^{\prime}}\|A\|_{\mathrm{L}^{p}(\mathbb{R}^{n+m};\mathrm{M}_{d}(\mathbb{C}))},$
@@
It is clear that the proofs of Propositions 8, 9 work also in the one-parameter setting.